12 HWK Assignment blem 12.10 ㎜ Review 1 Constants i Periodic Table PartA n ideal gas at O C consists of 1.0 x 1023 atoms. 11 J of ermal energy are added to the gas. What is the new temperature in C Express your answer in degrees Celsius to two significant figures 이 АЕф Submit X Incorrect; Try Again; 5 attempts remaining Provide Feedback
Answer #1

As we know that 1 mole of a gas consists of 6.023 x 10^23 atoms

So, 1.0 x 10^23 atoms = (1.0 x 10^23) / (6.023 x 10^23) = 0.166 mole

Assuming the process under constant volume

Cv for ideal gas =12.5 J/ (mol.K)

Amount of thermal energy added to the gas = 11.0 J

So -

μCv dt = 0.166 x 12.5 x dt = 11.0 J

=> dt = 11.0 / (0.166 x 12.5 ) = 5.30 K

Therefore, the new temperature of the gas = 0 + 5.30 = 5.30 deg C.

(Please note that we have considered the temperature 5.30 K is equal to 5.30 deg C, because when finding the difference in temperature, both units kelvin and deg. C have equal value.)